OpenStudy (anonymous):

Find a degree 3 polynomial with coefficient of x^3 equal to 1 and zeros -2, -2 i and 2 i

OpenStudy (campbell_st):

well the complex roots of x = -2i, 2i means the quadratic is \[x^2 = -4 ~~~~~~or~~~~~~~x^2 + 4 = 0\] the the general form is \[f(x) = a(x^2 + 4)(x + 2) \] you are told the coefficient of the leading term is 1 so a = 1 so you need to expand/distribute \[f(x) = (x^2 + 4)(x + 2)\] to find the polynomial.

OpenStudy (anonymous):

Thank you!